If (1/b+c) +(1/c+a)=(2/a+b)
find a^2+b^2
please give the solution.

Grade:12

3 Answers

bharat bajaj IIT Delhi
10 years ago
Take LCM ,
(2c+a+b)/(a+c)(b+c) = 2/a+b
Cross multiply:
a^2 + b^2 + 2ab + 2bc + 2ac = 2bc + 2ab + 2c^2 + 2ac
This gives,
a^2 + b^2 = 2. c^2
Thanks
Bharat Bajaj
askiitians faculty
IIT Delhi
bharat bajaj IIT Delhi
9 years ago
Take LCM ,
(2c+a+b)/(a+c)(b+c) = 2/a+b
Cross multiply:
a^2 + b^2 + 2ab + 2bc + 2ac = 2bc + 2ab + 2c^2 + 2ac
This gives,
a^2 + b^2 = 2. c^2


Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty
R1312467
9 years ago
Take LCM , (2c+a+b)/(a+c)(b+c) = 2/a+b Cross multiply: a^2 + b^2 + 2ab + 2bc + 2ac = 2bc + 2ab + 2c^2 + 2ac This gives, a^2 + b^2 = 2. c^2